OpenStudy (chosenmatt):

Mandana is ordering new tile for her kitchen floor. The area of the floor is 116.9 square feet, and she wants to order between 15% and 20% extra of the materials. Which of the following is a reasonable amount of tile for Mandana to order?

OpenStudy (chosenmatt):

@KamiBug

OpenStudy (chosenmatt):

@uri

OpenStudy (briensmarandache):

the ans btw 15% more= 116.9(0.15)+116.9 and 20% more= 116.9(0.2)+116.9

OpenStudy (chosenmatt):

133 square feet 138 square feet 141 square feet

OpenStudy (chosenmatt):

those are answer choices

OpenStudy (chosenmatt):

@KamiBug

OpenStudy (chosenmatt):

@KamiBug

OpenStudy (briensmarandache):

did you plug the number i gave you in the calculator?

OpenStudy (chosenmatt):

huh?

OpenStudy (chosenmatt):

im confuse

OpenStudy (briensmarandache):

15% more= 116.9(0.15)+116.9

OpenStudy (briensmarandache):

20% more= 116.9(0.2)+116.9

OpenStudy (briensmarandache):

percentage on top of anything is away the amount*% + amt because your adding, percentage away of anything is away the amount - amount*% because your adding,

OpenStudy (briensmarandache):

the last word of that statement should be subtracting

OpenStudy (chosenmatt):

@KamiBug

OpenStudy (briensmarandache):

did you get it?